OpenStudy (anonymous):

Simplify. Squareroot over full equation: b^(2n)/b^(-2n)

myininaya (myininaya):

\[\sqrt{b^{2n} \cdot b^{2n}}=\sqrt{b^{2n+2n}}=\sqrt{b^{4n}}=b^{\frac{4n}{2}}\]

myininaya (myininaya):

see if you can simplify 4/2